# Restoring the Forecasting Power of Google Trends with Statistical Preprocessing

### v1

@@ Abstract
Google Trends reports how frequently specific queries are searched on Google over time. It is widely used in research and industry to gain early insights into public interest. However, its data generation mechanism introduces missing values, sampling variability, noise, and trends. These issues arise from privacy thresholds mapping low search volumes to zeros, daily sampling variations causing discrepancies across historical downloads, and algorithm updates altering volume magnitudes over time. Data quality has recently deteriorated, with more zeros and noise, even for previously stable queries. We propose a comprehensive statistical methodology to preprocess Google Trends search information using hierarchical clustering, smoothing splines, and detrending. We validate our approach by forecasting U.S. influenza hospitalizations with a univariate ARIMAX model. Compared to omitting exogenous variables, our results show that raw Google Trends data degrades modeling performance, while preprocessed signals enhance forecast accuracy by 58% nationally and 24% at the state level.
Keywords: Google Trends, Preprocessing, Clustering, Denoising, Detrending, ARIMAX
@@ 1 Introduction
The digital footprint generated by billions of daily online searches offers a wealth of data, capturing public interests and behaviors across a broad range of topics. As the world’s leading search engine, Google plays a central role in this ecosystem, with Google Trends emerging as a powerful, public-facing tool for gaining early insights into evolving patterns. Over the past two decades, Google Trends has been successfully applied in various fields, including epidemiology to predict infectious diseases and outbreaks (Yang et al. 2015), macroeconomics to forecast unemployment rates (D’Amuri & Marcucci 2017), finance to analyze stock market trends (Preis et al. 2013), and marketing to study consumer behavior in fashion (Silva et al. 2019). Compared to traditional data sources such as surveys, these aggregated search volumes are low-cost, continuously updated, and can provide a more dynamic and accurate representation of how behaviors, opinions, and sentiments evolve over time (Stephens-Davidowitz 2017).
@@
1.1 Related Work
@@
Although there is extensive research on the use of Google Trends data for multiple applications, only a few recent analyses have focused on quantifying and addressing data quality issues. Some studies suggest removing all queries whose time series have missing values from the analysis (Cebrián & Domenech 2023, Liu et al. 2015), resulting in a potential loss of valuable information. To mitigate this, Neumann et al. (2023) recommend combining search phrases in the Google Trends tool using the Boolean operator OR (“+”), which allows the retrieval of aggregated search volumes. However, their approach does not describe how to effectively construct these combinations, leaving a gap in practical implementation.
@@
Sampling variability is another area of concern in Google Trends data. In Cebrián & Domenech (2023), the authors highlight that the lack of accuracy in Google Trends data originates from the sampling process, but do not propose specific methods to improve data quality. In contrast, Eichenauer et al. (2022) and Neumann et al. (2023) explain the Google Trends sampling mechanism, which introduces noise and fluctuations in the returned search volumes, and suggest averaging results from multiple downloads to reduce these variations. Expanding on this, Cebrián & Domenech (2024) model and simulate the data generation process to better understand sampling error, proposing a measure to determine the number of samples required for consistent search volumes. However, since Google servers cache query results and return identical values for repeated requests within a 24-hour period, averaging multiple samples may not be practical for real-time analysis.
@@
Several studies have applied temporal smoothing techniques to preprocess Google Trends data before integrating it into forecasting models. For instance, moving averages are employed to smooth time series of search volumes (Liu et al. 2015, Rabiolo et al. 2021, Wang et al. 2022, Lampos et al. 2021). However, this method can introduce phase shifts, particularly for weekly or monthly data where the peak is delayed after smoothing. Alternatively, Silva et al. (2019) utilize singular spectrum analysis to denoise search data, demonstrating its effectiveness in enhancing the performance of a neural network autoregression forecasting model. Similarly, Fenga (2020) combines a Seasonal AutoRegressive Integrated Moving Average (SARIMA) model with a wavelet denoising filter, showing that the addition of the filter improves forecasting accuracy compared to using SARIMA alone. While these studies effectively address noisy signals and underscore the benefits of denoising Google Trends data to enhance forecasts, their methodologies rely on the entire time series for smoothing, including future time periods. This makes them unsuitable for real-time forecasting, as future data is not available at the time of prediction. Retrospective application of these techniques could introduce future-looking information, potentially leading to unrealistic accuracy improvements not achievable in real-time settings.
@@
One of the earliest and most significant applications of Google Trends has been infectious disease forecasting in the United States (US). Inspired by its prior successes, influenza activity prediction at the national and state levels in the US serves as a case study to demonstrate the effectiveness of our Google Trends data preprocessing methodology. The first documented use of Google search data was presented by Ginsberg et al. (2009), who leveraged search volumes to forecast influenza outbreaks, leading to the development of Google Flu Trends (GFT) tool. Following the failure of GFT to provide reliable forecasts (Santillana et al. 2014), substantial research has been conducted on refining Google Trends-based predictions by integrating data from multiple sources and calibrating machine learning models (Yang et al. 2017, Kandula et al. 2019, Ning et al. 2019, Yang et al. 2021). In particular, Yang et al. (2015) improved the GFT framework with ARGO (AutoRegression with GOogle search data), an autoregressive model with LASSO regularization that accurately estimates influenza-like illness (ILI) rates based on historical flu activity and Google search data. Building on this, Ning et al. (2019) developed ARGO2 (2-step Augmented Regression with GOogle data), which refines ARGO by combining national and regional predictions, incorporating cross-regional correlations, and computing the final ILI estimates using the best linear predictor. The performance of these models shows the benefits of integrating Google Trends data in predictive models applied to influenza forecasting.
@@ 1.2 Our contributions
In this paper, we present a preprocessing methodology designed to enhance the quality of Google Trends data for real-time forecasting applications. Our approach addresses key challenges inherent in search data, such as missing values, sampling variability, noise, and trends, thereby improving its consistency, stability, and predictive utility. This reproducible framework refines Google Trends data into more reliable and informative features for prediction tasks.
We begin with a practical overview of Google Trends, highlighting its functionalities, challenges, and strategies for selecting relevant search queries. Our statistical methodology is then presented in three stages. (1) First, we employ hierarchical clustering to group similar keywords based on correlations between their time series. Combined search volumes of related terms overcome Google’s privacy thresholds, yielding non-zero values for initially sparse data, thus reducing missing values and stabilizing sampling variability. (2) Next, we apply smoothing splines iteratively to denoise the time series without introducing look-ahead bias, preserving the integrity necessary for real-time forecasting models. This step effectively reduces noise caused by daily sampling variations and short-term fluctuations in searches. (3) To further improve data quality, we use linear detrending, quadratic detrending, and differencing to eliminate long-term deterministic and stochastic trends, ensuring stationarity in the time series while preserving short-term dynamics. We then use correlations between the resulting predictors and the target variables to filter and retain only the most relevant search queries.
We validate our preprocessing framework through its application to predicting influenza hospitalizations in the United States using an AutoRegressive Integrated Moving Average with eXogenous variables (ARIMAX) model, incorporating Google Trends data as external predictors. Our results show that using preprocessed search data significantly enhances prediction accuracy compared to omitting external variables entirely, while using raw data degrades performance. Although our case study focuses on infectious disease forecasting, the proposed approach is designed as a general-purpose framework for preprocessing Google Trends data, making it suitable for a wide range of domains.
@@ 2 Data
Google Trends is the primary data source for this study, providing search volumes for flu-related terms. Available publicly at https://trends.google.com/trends/, it offers valuable insights into search behaviors but presents challenges due to its data generation mechanism. For an end-to-end application of our preprocessing methodology, we supplement this data with influenza hospital admission counts from the U.S. Centers for Disease Control and Prevention (CDC).
@@ 2.1 Description of Google Trends
Google Trends allows users to analyze search behavior for inputted keywords by location, time range, category and search type. The search volume is available for two types of keywords: “search terms” and “topics”. A search term can be a single word or a phrase, with returned volumes encompassing searches that include the words in any order, as well as with additional words before or after. Boolean logic can be used to retrieve specific phrases or exclude some terms (Trends Help 2025d). Topics are broader than search terms and typically yield higher search volumes. They are “generally considered to be more reliable for Google Trends data” (Google News Initiative 2025), as they account for all languages and spellings of a word. Moreover, results for topics are more refined and less noisy, as these are associated with a category (such as “Health” or “Finance”), which returns results in the desired field when a term has multiple meanings (Trends Help 2025c).
@@
For a specific search, there are four types of data representations: “interest over time”, “interest by subregion”,“related queries” and “related topics”. The “interest over time” feature returns a time series of search volumes that reflects the relative popularity of a keyword over time, scaled from 0 to 100, where 100 corresponds to the peak volume (Trends Help 2025a). This time series reveals trends, seasonal patterns, or spikes in interest, and can be obtained at various temporal resolutions, e.g., hourly, daily, weekly or monthly. Figure 1 displays a representative example of search volumes for the keyword “Cough”. “Interest by subregion” provides a geographical breakdown of search data, highlighting the areas where a subject is most popular. By explicitly selecting a location in the input box, the interest over time also provides the interest in that particular region.
@@
To further explore relevant search activity, the “related queries” and “related topics” functionalities help us identify a broader set of terms and topics connected to the initial search keyword. They can be selected as “rising”, that is, emerging terms that had a significant increase in search volume, or “top”, that is, most popular search terms in the period, location and category chosen (Trends Help 2025b). Hence, choosing different time periods, regions or categories will give different related searches. Section 2.3 provides a more detailed approach to obtaining relevant keywords.
@@ 2.2 Challenges of Google Trends Data
Google Trends assigns a value of “0” when the number of searches for a specific keyword falls below undisclosed privacy thresholds defined by Google, as illustrated in Figure 2(a). It indicates low search activity rather than a complete absence of searches. This can result in sparse data, particularly at the state or city level and for daily or weekly frequencies (Stephens-Davidowitz & Varian 2014). These zeros represent missing data that need to be addressed, as they mask low but potentially meaningful search activity.
@@
Noise, shown in Figure 2(b), is also inherent in the data generation process of Google Trends. Given the billions of Google searches conducted daily, Google Trends provides only a representative sample to ensure quick response times (Trends Help 2025a). This sample is a random approximation of the entire population of searches, and thus contains sampling noise. To optimize processing, Google servers cache data for 24 hours, meaning that repeated input searches within the same day yield identical results. However, at midnight UTC, the cache resets, and a new sample is drawn, leading to day-to-day variability across different downloads of search volumes for the same search (Choi & Varian 2012, Neumann et al. 2023), as illustrated in Figure 2(d). This variability arises from the randomness of the sampling-based data generation process, which follows a hypergeometric distribution (Bleher & Dimpfl 2022). Let $K$ and $k$ be the search volumes for a keyword in a population of size $N$ and a sample of size $n$, respectively. Since $K$ and $N$ are unobserved, the reported proportion $\frac{k}{n}$ is a noisy approximation of $\frac{K}{N}$. Google Trends scales search volumes from 0 to 100, effectively returning $100\times\frac{k}{n}/\frac{K}{N}$. As $K$ decreases, the variance of $\frac{k}{n}/\frac{K}{N}$, which reflects sampling variability, increases. This leads to noisier estimates, especially for small regions and infrequent searches.
As terms become more or less popular with rising or declining public interest, search volumes may exhibit a trend, as displayed in Figure 2(c). It corresponds to a gradual change in the data over time, which can mask underlying patterns. In the context of Google Trends, this phenomenon can have multiple causes, including changes in vocabulary with new or obsolete terms, more or less general interest in a subject, or algorithm updates.
@@
In fact, data quality could change as Google Trends’ algorithm undergoes updates. Cebrián & Domenech (2024) mention that there has been notable changes in Google Trends’ algorithms over the years. According to the Google Trends website, improvements to geographical assignment and data collection systems were implemented in 2011, 2016, and 2022, as shown in SI Figure S1. These updates can lead to inconsistencies. For example, Myburgh (2022) reports that changes to Google’s sampling strategy in January 2022 resulted in higher search volumes compared to earlier data. Consequently, data after January 2022 is not directly comparable to prior periods. Therefore, long-term trends may reflect algorithmic changes rather than genuine shifts in search behavior, posing challenges for accurate forecasting.
@@
2.3 Identifying Flu-related Terms from Google Trends
In this study, we analyze weekly Google Trends’ “interest over time” data for the United States at both national and state levels since 2004. Our initial set of 161 flu-related terms and topics is based on the work of Yang et al. (2015), identified through Google Correlate in 2009 and 2010, but the service was discontinued in 2019. Therefore, to expand our pool of keywords, we retrieve queries related to “flu”. For long-term seasonal analysis, “top” queries are preferred over “rising” ones, as they exhibit more consistent search patterns in the given time frame. We choose the “Cold and Flu” category to limit results to the area of infectious diseases, and select the entire period from 2004 onward to exclude event-specific queries (e.g., “flu 2024”). Keywords are tailored to each state by setting the corresponding location, which also restricts results to English (or the predominant local language), as opposed to collecting worldwide data. In summary, category, time period, and location serve as filters to refine the dataset. This approach yields time series of weekly search volumes for approximately 500 terms per state, covering symptoms, treatments, medicine, or types of flu.
2.4 Influenza Hospitalizations from CDC
Every year since 2013, the US Centers for Disease Control and Prevention organizes a “Predict the Influenza Season Challenge”, also called “FluSight”, which asks participating teams to forecast weekly influenza-related metrics in real-time. Initially requiring forecasts of outpatient influenza-like-illness (ILI) rates, the target variable shifted in 2021 to flu-related hospitalizations up to 4 weeks in the future (Mathis et al. 2024, Centers for Disease Control and Prevention 2025a). Several prediction methods are employed, generally together in an ensemble model, including statistical models like ARIMA (Kandula & Shaman 2019, Meyer et al. 2025), machine learning models such as gradient boosting (Meyer et al. 2025, Ray et al. 2024), mechanistic models with compartments (Aawar & Srivastava 2022), deep learning models like neural networks (Aiken et al. 2021). We use the data from this challenge to retrospectively evaluate the performance of using preprocessed Google Trends search volumes in predictive models.
Reporting of hospitalization data became mandatory on February 2, 2022 (Mathis et al. 2024), and admission counts, collected at both national and state levels in the United States by the National Healthcare Safety Network (NHSN) (Centers for Disease Control and Prevention 2025d) are publicly available in the GitHub repository from CDC Epidemic Prediction Initiative (2025). Influenza-related hospitalization data is released weekly with a one-week delay, while Google Trends data is available in real time, offering a timely informational advantage. In this study, we leverage search data as exogenous variables to perform 1-week-ahead prediction of flu hospitalizations.
To extend the training period for modeling, we follow the methodology of Meyer et al. (2025) to augment the dataset using ILINet data (Centers for Disease Control and Prevention 2025c) and FluSurv-Net data (Centers for Disease Control and Prevention 2025b) with transfer learning. This augmentation provides a continuous time series of hospitalization data starting from 2012.
3 Methods

### v2

@@ Abstract
Google Trends reports how frequently specific queries are searched on Google over time. It is widely used in research and industry to gain early insights into public interest. However, its data generation mechanism introduces missing values, sampling variability, noise, and trends. These issues arise from privacy thresholds mapping low search volumes to zeros, daily sampling variations causing discrepancies across historical downloads, and algorithm updates altering volume magnitudes over time. Data quality has recently deteriorated, with more zeros and noise, even for previously stable queries. We propose a comprehensive statistical methodology to preprocess Google Trends search information using hierarchical clustering, smoothing splines, and detrending. We validate our approach by forecasting U.S. influenza hospitalizations up to three weeks ahead with several statistical and machine learning models. Compared to omitting exogenous variables, our results show that preprocessed signals enhance forecast accuracy, while raw Google Trends data often degrades performance in statistical models.
keywords:
Google Trends , Preprocessing , Clustering , Denoising , Detrending , Forecasting
journal: International Journal of Forecasting
[…]
@@ 1 Introduction
The digital footprint generated by billions of daily online searches offers a wealth of data, capturing public interests and behaviors across a broad range of topics. As the world’s leading search engine, Google plays a central role in this ecosystem, with Google Trends emerging as a powerful, public-facing tool for gaining early insights into evolving patterns. Over the past two decades, Google Trends has been successfully applied in various fields, including epidemiology to predict infectious diseases and outbreaks (Yang et al., 2015), macroeconomics to forecast unemployment rates (D’Amuri and Marcucci, 2017), finance to analyze stock market trends (Preis et al., 2013), and marketing to study consumer behavior in fashion (Silva et al., 2019). One of the earliest and pioneering works in this area, (Choi and Varian, 2012), showed that Google Trends data can enhance real-time estimation of economic indicators, with their 2009 technical report (Choi and Varian, 2009) illustrating how search behavior can help “predict the present”. Compared to traditional data sources such as surveys, these aggregated search volumes are low-cost, continuously updated, and can provide a more dynamic and accurate representation of how behaviors, opinions, and sentiments evolve over time (Stephens-Davidowitz, 2017).
@@
1.1 Outline of the paper
In this paper, we present a preprocessing methodology designed to enhance the quality of Google Trends data for real-time forecasting applications. Our approach addresses key challenges inherent in search data, such as missing values, sampling variability, noise, and trends, thereby improving its consistency, stability, and predictive utility. We begin with a literature review and a dedicated data section providing a practical overview of Google Trends, highlighting its functionalities, challenges, and strategies for selecting relevant search queries. Our statistical methodology is then presented in three stages as illustrated in Figure 1. (1) First, we employ hierarchical clustering to group similar keywords based on correlations between their time series. Combined search volumes of related terms overcome Google’s privacy thresholds, yielding non-zero values for initially sparse data, thus reducing missing values and stabilizing sampling variability. (2) Next, we apply smoothing splines iteratively to denoise the time series without introducing look-ahead bias, preserving the integrity necessary for real-time forecasting models. This step effectively reduces noise caused by daily sampling variations and short-term fluctuations in searches. (3) To further improve data quality, we use linear detrending, quadratic detrending, and differencing to eliminate long-term deterministic and stochastic trends, ensuring stationarity in the time series while preserving short-term dynamics. We then use correlations between the resulting predictors and the target variables to filter and retain only the most relevant search queries. We validate our preprocessing framework through its application to predicting influenza hospitalizations in the United States using three statistical and two machine learning forecasting models, incorporating Google Trends data as external predictors.
Figure 1: Flowchart of the three-step preprocessing methodology for Google Trends data. Orange circles represent the search volumes retrieved from the Google Trends interface, blue circles represent the outputs produced by each preprocessing step.
2 Related Work
@@
Although there is extensive research on the use of Google Trends data for multiple applications, only a few recent analyses have focused on quantifying and addressing data quality issues. Some studies suggest removing all queries whose time series have missing values from the analysis (Cebrián and Domenech, 2023; Liu et al., 2015), resulting in a potential loss of valuable information. To mitigate this, Neumann et al. (2023) recommend combining search phrases in the Google Trends tool using the Boolean operator OR (“+”), which allows the retrieval of aggregated search volumes. However, their approach does not describe how to effectively construct these combinations, leaving a gap in practical implementation.
@@
Sampling variability is another area of concern in Google Trends data. In Cebrián and Domenech (2023), the authors highlight that the lack of accuracy in Google Trends data originates from the sampling process, but do not propose specific methods to improve data quality. In contrast, Eichenauer et al. (2022) and Neumann et al. (2023) explain the Google Trends sampling mechanism, which introduces noise and fluctuations in the returned search volumes, and suggest averaging results from multiple downloads to reduce these variations. Expanding on this, Cebrián and Domenech (2024) model and simulate the data generation process to better understand sampling error, proposing a measure to determine the number of samples required for consistent search volumes. However, since Google servers cache query results and return identical values for repeated requests within a 24-hour period, averaging multiple samples may not be practical for real-time analysis.
@@
Several studies have applied temporal smoothing techniques to preprocess Google Trends data before integrating it into forecasting models. For instance, moving averages are employed to smooth time series of search volumes (Liu et al., 2015; Rabiolo et al., 2021; Wang et al., 2022; Lampos et al., 2021). However, this method can introduce phase shifts, particularly for weekly or monthly data where the peak is delayed after smoothing. Alternatively, Silva et al. (2019) utilize singular spectrum analysis to denoise search data, demonstrating its effectiveness in enhancing the performance of a neural network autoregression forecasting model. Similarly, Fenga (2020) combines a Seasonal AutoRegressive Integrated Moving Average (SARIMA) model with a wavelet denoising filter, showing that the addition of the filter improves forecasting accuracy compared to using SARIMA alone. While these studies effectively address noisy signals and underscore the benefits of denoising Google Trends data to enhance forecasts, their methodologies rely on the entire time series for smoothing, including future time periods. This makes them unsuitable for real-time forecasting, as future data is not available at the time of prediction. Retrospective application of these techniques could introduce future-looking information, potentially leading to unrealistic accuracy improvements not achievable in real-time settings.
@@
One of the earliest and most significant applications of Google Trends has been infectious disease forecasting in the United States (US). Influenza prediction is a well-studied application of Google Trends data and therefore serves as the case study to demonstrate the effectiveness of our preprocessing methodology. The first documented use of Google search data was presented by Ginsberg et al. (2009), who leveraged search volumes to forecast influenza outbreaks, leading to the development of Google Flu Trends (GFT) tool. Following the failure of GFT to provide reliable forecasts (Santillana et al., 2014), substantial research has been conducted on refining Google Trends-based predictions by integrating data from multiple sources and calibrating machine learning models (Yang et al., 2017; Kandula et al., 2019; Ning et al., 2019; Yang et al., 2021). In particular, Yang et al. (2015) improved the GFT framework with ARGO (AutoRegression with GOogle search data), an autoregressive model that selects the most relevant exogenous predictors to accurately estimate influenza-like illness (ILI) rates based on historical flu activity and Google search data. Building on this, Ning et al. (2019) developed ARGO2 (2-step Augmented Regression with GOogle data), which refines ARGO by combining national and regional predictions, incorporating cross-regional correlations, and computing the final ILI estimates using the best linear predictor. The performance of these models shows the benefits of integrating Google Trends data in predictive models applied to influenza forecasting.
@@ 2.1 Our Contributions
Our contributions are threefold. (1) We present a general preprocessing methodology that enhances the quality of Google Trends data to improve real-time forecasting performance. (2) Through the proposed approach, we address multiple challenges inherent in search data, including missing values, sampling variability, noise, and long-term trends. The framework is fully reproducible and designed for practical implementation. It consists of a three-step procedure that transforms raw search volumes into reliable and informative features while preserving the integrity required for real-time settings. (3) We apply our methods to influenza hospitalization forecasting across diverse statistical and machine learning models. Our results show that using preprocessed search data significantly increases accuracy compared to omitting external variables, while using raw data degrades performance in statistical models. Although our case study focuses on infectious disease, the proposed approach is broadly applicable to other domains where Google Trends data can be used as an exogenous indicator.
@@ 3 Data
Google Trends is the primary data source for this study, providing search volumes for flu-related terms. Available publicly at https://trends.google.com/trends/, it offers valuable insights into search behaviors but presents challenges due to its data generation mechanism.
@@ 3.1 Description of Google Trends
Google Trends allows users to analyze search behavior for inputted keywords by location, time range, category and search type. The search volume is available for two types of keywords: “search terms” and “topics”. A search term can be a single word or a phrase, with returned volumes encompassing searches that include the words in any order, as well as with additional words before or after. Boolean logic can be used to retrieve specific phrases or exclude some terms (Trends Help, 2025d). Topics are broader than search terms and typically yield higher search volumes. They are “generally considered to be more reliable for Google Trends data” (Google News Initiative, 2025), as they account for all languages and spellings of a word. Moreover, results for topics are more refined and less noisy, as these are associated with a category (such as “Health” or “Finance”), which returns results in the desired field when a term has multiple meanings (Trends Help, 2025c).
@@
For a specific search, there are four types of data representations: “interest over time”, “interest by subregion”,“related queries” and “related topics”. The “interest over time” feature returns a time series of search volumes that reflects the relative popularity of a keyword over time, scaled from 0 to 100, where 100 corresponds to the peak volume (Trends Help, 2025a). This time series reveals trends, seasonal patterns, or spikes in interest, and can be obtained at various temporal resolutions, e.g., hourly, daily, weekly or monthly. Figure 2 displays a representative example of search volumes for the keyword “Cough”. “Interest by subregion” provides a geographical breakdown of search data, highlighting the areas where a subject is most popular. By explicitly selecting a location in the input box, the interest over time also provides the interest in that particular region.
@@
To further explore relevant search activity, the “related queries” and “related topics” functionalities help us identify a broader set of terms and topics connected to the initial search keyword. They can be selected as “rising”, that is, emerging terms that had a significant increase in search volume, or “top”, that is, most popular search terms in the period, location and category chosen (Trends Help, 2025b). Hence, choosing different time periods, regions or categories will give different related searches. Section 6.1 provides a more detailed approach to obtaining relevant keywords.
@@ 3.2 Challenges of Google Trends Data
Google Trends assigns a value of “0” when the number of searches for a specific keyword falls below undisclosed privacy thresholds defined by Google, as illustrated in Figure 3(a). It indicates low search activity rather than a complete absence of searches. This can result in sparse data, particularly at the state or city level and for daily or weekly frequencies (Stephens-Davidowitz and Varian, 2014). These zeros represent missing data that need to be addressed, as they mask low but potentially meaningful search activity.
@@
Noise, shown in Figure 3(b), is also inherent in the data generation process of Google Trends. Given the billions of Google searches conducted daily, Google Trends provides only a representative sample to ensure quick response times (Trends Help, 2025a). This sample is a random approximation of the entire population of searches, and thus contains sampling noise. To optimize processing, Google servers cache data for 24 hours, meaning that repeated input searches within the same day yield identical results. However, at midnight UTC, the cache resets, and a new sample is drawn, leading to day-to-day variability across different downloads of search volumes for the same search (Choi and Varian, 2012; Neumann et al., 2023), as illustrated in Figure 3(d). This variability arises from the randomness of the sampling-based data generation process, which follows a hypergeometric distribution (Bleher and Dimpfl, 2022). Let $K$ and $k$ be the search volumes for a keyword in a population of size $N$ and a sample of size $n$, respectively. Since $K$ and $N$ are unobserved, the reported proportion $\frac{k}{n}$ is a noisy approximation of $\frac{K}{N}$. Google Trends scales search volumes from 0 to 100, effectively returning $100\times\frac{k}{n}/\frac{K}{N}$. As $K$ decreases, the variance of $\frac{k}{n}/\frac{K}{N}$, which reflects sampling variability, increases. This leads to noisier estimates, especially for small regions and infrequent searches.
As terms become more or less popular with rising or declining public interest, search volumes may exhibit a trend, as displayed in Figure 3(c). These trends represent gradual changes over time that can mask underlying patterns. In the context of Google Trends, this phenomenon can have multiple causes, including changes in vocabulary with new or obsolete terms, shifts in general public interest in a subject, or the platform’s dynamic normalization. Specifically, Google Trends scales its “interest over time” index from 0 to 100 within the selected time period and region. When the temporal or spatial window changes, the entire series is rescaled so that the new maximum equals 100, and all other values are adjusted proportionally. This can introduce artificial level shifts, low-frequency drifts, and apparent structural breaks that resemble long-term trends but do not reflect actual changes in search behavior. Such spurious trends can bias correlations and obscure long-term relationships with the target variable.
@@
Additionally, data quality could change as Google Trends’ algorithm undergoes updates. Cebrián and Domenech (2024) mention that the platform has undergone notable algorithmic updates over the years. According to the Google Trends website, improvements to geographical assignment and data collection systems were implemented in 2011, 2016, and 2022, as shown in SI Figure S1. These updates can lead to inconsistencies. For example, Myburgh (2022) reports that changes to Google’s sampling strategy in January 2022 resulted in higher search volumes compared to earlier data. Consequently, data after January 2022 is not directly comparable to prior periods. Therefore, long-term trends may also reflect algorithmic changes rather than genuine shifts in search behavior, posing challenges for accurate forecasting.
@@
4 Methodology